OpenStudy (anonymous):

Find the total area of the regular pyramid. Answer in the form of: (XX(Sqrt) XX + X (Sqrt) X)

OpenStudy (anonymous):

https://media.glynlyon.com/g_geo_2012/8/groupi78.gif

OpenStudy (anonymous):

Surface area? @AtomicFishbowl

OpenStudy (anonymous):

The total surface area, I believe so.

OpenStudy (aivantettet26):

nps^2 all over 4tan(180/n)

OpenStudy (anonymous):

What is nps? And if I work out the problem that way will you fall into the form I need it to be in?

OpenStudy (anonymous):

I also don't have a scientific calculator.

OpenStudy (anonymous):

Is the base a regular hexagon? @AtomicFishbowl

OpenStudy (aivantettet26):

n is the number of sides p is the number of polygons enclosing the polyhedron e is the number of edges

OpenStudy (anonymous):

Yes, I believe so.

OpenStudy (anonymous):

I don't understand what p is.

OpenStudy (aivantettet26):

do you know the final answer?

OpenStudy (anonymous):

I have no idea...

OpenStudy (anonymous):

Will you tell me how to get the final answer in that form please?

OpenStudy (anonymous):

@AtomicFishbowl Notice that that you can divide this hexagon-based pyramid in to 6 triangle based pyramid. Do you see?

OpenStudy (aivantettet26):

yes i will :)

OpenStudy (anonymous):

Yes I thought about using the Pythagorean theorem but I don't think I have enough information for that

OpenStudy (anonymous):

And even if I did I wouldn't know what to do with it.

OpenStudy (anonymous):

OK now. Notice that because that the base is a regular hexagon, each triangle based pyramid is the SAME and each has base that is an equilateral triangle with length 6. Are you with me? @AtomicFishbowl

OpenStudy (aivantettet26):

my answer is 654.72 units squared

OpenStudy (anonymous):

@aivantettet26 But it has to be in the form of (XX(Sqrt)XX + X (Sqrt) X) OHH! @genius12 I didn't put together that it was an equilateral triangle. I'm with you, go on.

OpenStudy (aivantettet26):

i mean that's my answer not really the final answer. im waiting for the answer of sir genius12 :)

OpenStudy (anonymous):

Lol I just remembered this is surface area. I was thinking of Volume the whole time rofl...Anyway, I'll do the calculations right now and provide the steps if you need it.

OpenStudy (anonymous):

Okay, whichever one is fine! I'd appreciate the answer but if you can explain your steps I'd like that too.

OpenStudy (anonymous):

Anyway, notice that since the base is a regular hexagon, we can divide the base in two six equilateral triangle each with side length 6. The height of each this equilateral triangles is \(\bf \frac{\sqrt{3}}{2}a\) where a = 6 hence the height is \(\bf 3\sqrt{3}\) of each equilateral triangle. Employing the pythagorean theorem, we can find the height of the "outer isosceles triangles" with base 6 and the height of these outer triangles will be:\[\bf (8)^2+(3\sqrt{3})^2=64+27=91 \implies h = \sqrt{91}\]Now we can calculate the area of each of these outer isosceles triangles with base 6 and height \(\bf \sqrt{91}\) with the triangle area formula:\[\bf A=\frac{ 6 \times \sqrt{91} }{ 2 }=3\sqrt{91}\]Now since there 6 of these triangles, the total area of the triangles will be:\[\bf 6 \times 3\sqrt{91}=18\sqrt{91}\]Now we calculate the area of the base which is a hexagon that can be divided once again to those 6 same equilateral triangles with height \(\bf 3\sqrt{3}\). The area of each of these equilateral triangles is given by the same triangle area formula:\[\bf \frac{6 \times 3\sqrt{3}}{2}=9\sqrt{3}\]Now there is 6 of them so the total area of the hexagon becomes:\[\bf 6 \times 9\sqrt{3}=54\sqrt{3}\]Now we add the two final values we got and so the total surface area is:\[\bf 54\sqrt{3}+18\sqrt{91}=265.24 \ units^2\]

OpenStudy (anonymous):

@AtomicFishbowl @aivantettet26

OpenStudy (aivantettet26):

very nice !